# AURA: A Reinforcement Learning Framework for AI-Driven Adaptive Conversational Surveys

### v1

@@ 3.2 Data Collection and Preprocessing
To initialize AURA’s reinforcement-learning policy and establish empirical foundations for engagement measurement, we used campus-climate survey conversations collected from a pilot deployment of TigerGPT (Tang et al., 2025). TigerGPT is a semi-adaptive survey chatbot that conducts natural, open-ended dialogues but does not employ reinforcement learning. The resulting dataset provides authentic examples of engagement in conversational surveys and enables offline estimation of action effectiveness across different engagement states.
@@ 3.3.1 Foundations and Operationalization
Survey methodologists emphasize that response quality in open-ended questions is inherently multidimensional: no single indicator fully captures informativeness, engagement, or authenticity (Dillman et al., 2014; Tourangeau et al., 2000). Traditional assessment relies on post-hoc human coding, which cannot support real-time adaptive systems. AURA operationalizes four quality dimensions through automated linguistic analysis, enabling exchange-by-exchange measurement during conversation.
@@
Response length provides a first-order indicator of respondent investment and information content (Dillman et al., 2014; Holland and Christian, 2009). In traditional web surveys, detailed answers ($\approx$50 words) are rated as substantive, but chatbot conversation exchanges are naturally shorter because users elaborate across multiple exchanges. In our dataset ($N=467$), the median response length was 10 words and the 75th percentile (Q75) was 29 words, providing an empirically grounded scale for this dialogue context. Length is therefore normalized as
@@
First-person pronouns indicate self-focus and willingness to share personal experiences (Pennebaker et al., 2003; Tausczik and Pennebaker, 2010). In survey contexts, self-disclosure enriches qualitative feedback because responses grounded in personal experience (e.g., “I felt excluded when…”) reveal deeper insight than abstract statements (e.g., “Some students experience exclusion”). Because our responses are brief (median = 10 words), we use absolute rather than ratio-based pronoun counts to avoid mechanical dependence on response length: as text length increases, ratio scores decline even when disclosure remains constant, producing the strong Length–Disclosure correlation ($r=0.856$; Table 2). Absolute counts, capped at the 75th-percentile value of three pronouns, measure disclosure directly while retaining discriminative range. Self-disclosure is therefore normalized as
@@
Emotional language reflects psychological investment and topic salience. According to Russell’s (1980) circumplex model of affect, emotion varies along valence (positive–negative) and arousal (intensity). In surveys, highly affective language—positive or negative—indicates engagement (Porter and Whitcomb, 2005). To measure this automatically, we apply VADER sentiment analysis (Hutto and Gilbert, 2014) and use the magnitude of its compound score as a proxy for emotional intensity, independent of polarity. Although VADER estimates valence rather than arousal directly, research on affective norms shows that extreme valence correlates moderately with arousal ($r\approx 0.40$; Warriner et al. 2013), making the absolute compound value a practical indicator of engagement strength in brief conversational responses. Emotional intensity is therefore normalized as:
@@
Concrete, detailed responses convey richer and more actionable information than vague generalities. Episodic memory research shows that specific recollection involves three components—entities (who/what), temporal references (when), and spatial references (where) (Levine et al., 2002). Authentic accounts contain more such contextual details than fabricated ones ($d=0.72$; Newman et al. 2003). Unlike word-level concreteness metrics (Brysbaert et al., 2014), AURA detects the presence of these episodic components as binary indicators of specificity.
Each response is classified by GPT-4o for presence (1) or absence (0) of entities, temporal, and spatial details, following Levine et al.’s framework. Manual inspection of 50 samples confirmed consistent classification accuracy, after which the protocol was applied to the full corpus:
@@ 3.4.1 Taxonomy Development: A Theory-Informed, Data-Driven Approach
Action taxonomy development followed a hybrid methodology balancing empirical observation with theoretical grounding. Analysis of 467 chatbot questions across 96 TigerGPT conversations (Section 3.2) revealed recurring functional patterns: some prompts requested concrete examples, others sought elaboration, introduced new topics, or offered acknowledgment. These patterns were consolidated into five communicative categories and validated against established communication frameworks to ensure conceptual coherence (Li et al., 2016; Su et al., 2021).
@@ 3.4.2 Action Classification Methodology
Primary intent was defined as the question’s dominant illocutionary force (Searle, 1976)—the central action the question seeks to elicit. Model outputs included structured JSON records with confidence scores and brief reasoning to support quality control and flag ambiguous cases.
@@
    Note: CAT = Communication Accommodation Theory (Giles et al., 1991).
@@
    Speech act categories follow Searle (1976)’s taxonomy of illocutionary forces.
@@ 3.4.3 Action Type Definitions and Theoretical Grounding
Table 4 presents the five action types with their theoretical foundations and representative examples from the TigerGPT corpus. Each action corresponds to distinct constructs from Speech Act Theory (Searle, 1976) and Communication Accommodation Theory (CAT) (Giles et al., 1991), ensuring that AURA’s taxonomy captures functionally meaningful communicative strategies rather than superficial lexical patterns.
@@
Topic probe introduces new experiential dimensions, shifting attention to unexplored areas. These prompts function at the intersection of commissive and directive acts (Searle, 1976), committing the system to expanding the dialogue scope. They reflect CAT’s divergence strategy (Giles et al., 1991), ensuring broad coverage of campus climate dimensions (academic, social, financial, inclusivity).
Validation acknowledges user contributions without requesting new information. As expressive acts, validation supports rapport and psychological safety through maintenance strategies (Giles et al., 1991). Although rare in prior data ($n=4$, 0.9%), validation was retained because evidence shows it facilitates deeper disclosure on sensitive topics (Wheeless and Grotz, 1977; Melville and Hincks, 2016).
Continuation invites further input without specifying direction, granting users conversational autonomy. Also infrequent ($n=2$, 0.4%), continuation serves as a flow-maintaining strategy that may prove effective for users preferring open-ended engagement (Kocaballi et al., 2019).
@@ 3.6.2 Online Learning: Within-Session Adaptation
where $\alpha=0.3$ is the learning rate. This standard temporal-difference update (Sutton et al., 1998) occurs at exchange $t$ after observing the outcome of action $a_{t-1}$ selected in state $s_{t-1}$ at the previous exchange, increasing EV when the observed reward exceeds the prior estimate and decreasing it when the action underperforms.
@@ 6 Discussion and Future Work
The findings demonstrate that reinforcement learning can meaningfully enhance conversational survey systems by enabling real-time adaptation within individual sessions. AURA’s reinforcement-driven questioning strategy achieved consistent quality improvements compared with non-adaptive baselines, confirming that even short survey interactions benefit from dynamic policy updates guided by engagement feedback. The observed +0.12 mean gain and medium-to-large effect size ($d=0.66$) show that within-session learning is both statistically and practically significant.

### v2

@@ 3.2 Data Collection and Preprocessing
To initialize AURA’s reinforcement-learning policy and establish empirical foundations for engagement measurement, we used campus-climate survey conversations collected from a pilot deployment of TigerGPT (tang2025tigergpt). TigerGPT is a semi-adaptive survey chatbot that conducts natural, open-ended dialogues but does not employ reinforcement learning. The resulting dataset provides authentic examples of engagement in conversational surveys and enables offline estimation of action effectiveness across different engagement states.
@@ 3.3.1 Foundations and Operationalization
Survey methodologists emphasize that response quality in open-ended questions is inherently multidimensional: no single indicator fully captures informativeness, engagement, or authenticity (dillman2014internet; tourangeau2000psychology). Traditional assessment relies on post-hoc human coding, which cannot support real-time adaptive systems. AURA operationalizes four quality dimensions through automated linguistic analysis, enabling exchange-by-exchange measurement during conversation.
@@
Response length provides a first-order indicator of respondent investment and information content (dillman2014internet; holland2009measuring). In traditional web surveys, detailed answers ($\approx$50 words) are rated as substantive, but chatbot conversation exchanges are naturally shorter because users elaborate across multiple exchanges. In our dataset ($N=467$), the median response length was 10 words and the 75th percentile (Q75) was 29 words, providing an empirically grounded scale for this dialogue context. Length is therefore normalized as
@@
First-person pronouns indicate self-focus and willingness to share personal experiences (pennebaker2003linguistic; tausczik2010psychological). In survey contexts, self-disclosure enriches qualitative feedback because responses grounded in personal experience (e.g., “I felt excluded when…”) reveal deeper insight than abstract statements (e.g., “Some students experience exclusion”). Because our responses are brief (median = 10 words), we use absolute rather than ratio-based pronoun counts to avoid mechanical dependence on response length: as text length increases, ratio scores decline even when disclosure remains constant, producing the strong Length–Disclosure correlation ($r=0.856$; Table 2). Absolute counts, capped at the 75th-percentile value of three pronouns, measure disclosure directly while retaining discriminative range. Self-disclosure is therefore normalized as
@@
Emotional language reflects psychological investment and topic salience. According to Russell’s (russell1980circumplex) circumplex model of affect, emotion varies along valence (positive–negative) and arousal (intensity). In surveys, highly affective language—positive or negative—indicates engagement (porter2005mail). To measure this automatically, we apply VADER sentiment analysis (hutto2014vader) and use the magnitude of its compound score as a proxy for emotional intensity, independent of polarity. Although VADER estimates valence rather than arousal directly, research on affective norms shows that extreme valence correlates moderately with arousal ($r\approx 0.40$; warriner2013norms), making the absolute compound value a practical indicator of engagement strength in brief conversational responses. Emotional intensity is therefore normalized as:
@@
Concrete, detailed responses convey richer and more actionable information than vague generalities. Episodic memory research shows that specific recollection involves three components—entities (who/what), temporal references (when), and spatial references (where) (levine2002autobiographical). Authentic accounts contain more such contextual details than fabricated ones ($d=0.72$; newman2003lying). Unlike word-level concreteness metrics (brysbaert2014concreteness), AURA detects the presence of these episodic components as binary indicators of specificity.
Each response is classified by GPT-4o for presence (1) or absence (0) of entities, temporal, and spatial details, following levine2002autobiographical’s framework. Manual inspection of 50 samples confirmed consistent classification accuracy, after which the protocol was applied to the full corpus:
@@ 3.4.1 Taxonomy Development: A Theory-Informed, Data-Driven Approach
Action taxonomy development followed a hybrid methodology balancing empirical observation with theoretical grounding. Analysis of 467 chatbot questions across 96 TigerGPT conversations (Section 3.2) revealed recurring functional patterns: some prompts requested concrete examples, others sought elaboration, introduced new topics, or offered acknowledgment. These patterns were consolidated into five communicative categories and validated against established communication frameworks to ensure conceptual coherence (li2016deep; su2021dynamic).
@@ 3.4.2 Action Classification Methodology
Primary intent was defined as the question’s dominant illocutionary force (searle1976classification)—the central action the question seeks to elicit. Model outputs included structured JSON records with confidence scores and brief reasoning to support quality control and flag ambiguous cases.
@@
    Note: CAT = Communication Accommodation Theory (giles1991contexts).
@@
    Speech act categories follow searle1976classification’s taxonomy of illocutionary forces.
@@ 3.4.3 Action Type Definitions and Theoretical Grounding
Table 4 presents the five action types with their theoretical foundations and representative examples from the TigerGPT corpus. Each action corresponds to distinct constructs from Speech Act Theory (searle1976classification) and Communication Accommodation Theory (CAT) (giles1991contexts), ensuring that AURA’s taxonomy captures functionally meaningful communicative strategies rather than superficial lexical patterns.
@@
Topic probe introduces new experiential dimensions, shifting attention to unexplored areas. These prompts function at the intersection of commissive and directive acts (searle1976classification), committing the system to expanding the dialogue scope. They reflect CAT’s divergence strategy (giles1991contexts), ensuring broad coverage of campus climate dimensions (academic, social, financial, inclusivity).
Validation acknowledges user contributions without requesting new information. As expressive acts, validation supports rapport and psychological safety through maintenance strategies (giles1991contexts). Although rare in prior data ($n=4$, 0.9%), validation was retained because evidence shows it facilitates deeper disclosure on sensitive topics (wheeless1977measurement; melville2016conducting).
Continuation invites further input without specifying direction, granting users conversational autonomy. Also infrequent ($n=2$, 0.4%), continuation serves as a flow-maintaining strategy that may prove effective for users preferring open-ended engagement (kocaballi2019personalization).
@@ 3.6.2 Online Learning: Within-Session Adaptation
where $\alpha=0.3$ is the learning rate. This standard temporal-difference update (sutton1998reinforcement) occurs at exchange $t$ after observing the outcome of action $a_{t-1}$ selected in state $s_{t-1}$ at the previous exchange, increasing EV when the observed reward exceeds the prior estimate and decreasing it when the action underperforms.
@@ 6 Discussion and Future Work
The findings demonstrate that reinforcement learning can meaningfully enhance conversational survey systems by enabling real-time adaptation within individual sessions. AURA’s reinforcement-driven questioning strategy achieved consistent quality improvements compared with non-adaptive baselines, confirming that even short survey interactions benefit from dynamic policy updates guided by engagement feedback. The observed +0.076 mean gain and medium-to-large effect size ($d=0.66$) show that within-session learning is both statistically and practically significant.